A car travels 50 m at a rate of 10m/s,S. It then travels 200m at a rate of 20m/s, N. What is its average speed? Its average velocity??
Answer by mananth(16946)

50 m @ 10m/s
time = 50/10 = 5 seconds
200m @ 20m/s
time = 200/20 = 10 seconds
...
Total time taken = 20 +4 = 25 seconds
Total distance = 250m
...
Average Speed = 250/25 = 10 m/second.